2.15. Convert the circui indicated. ts as Electrical Circuit Concepts 101 (a) Use Z for all elements. a a o 9Ω bo (b) Use Y for all elements. a a o
Answer #1

@) Both ar and jobs are in Series į » Lucem ۱ Combination is z = 9+3162, Convension: R and jxq are in Combination is z = R(5x22+ > R²+x? 7 R²+x2 = 5 - 12 pot = 16 R²+x22 Equating o and ③ we get XY - GDR X = 0.5638 Substitusing X value in equation weo > &as islas as and jibs are in Paraines, Combination is Y - 9+s. - Y b- Conversion: G and i Be are in Combination is SeriesEquating equations 1 and ② we set by cope > Be - Gala > 0 · Be = 0.563 G Substituting Be in equation is we get G+(0.56365. G(